# DeciMamba: Exploring the Length Extrapolation Potential of Mamba

### v1

@@ Abstract
Long-range sequence processing poses a significant challenge for Transformers due to their quadratic complexity in input length. A promising alternative is Mamba, which demonstrates high performance and achieves Transformer-level capabilities while requiring substantially fewer computational resources. In this paper we explore the length-generalization capabilities of Mamba, which we find to be relatively limited. Through a series of visualizations and analyses we identify that the limitations arise from a restricted effective receptive field, dictated by the sequence length used during training. To address this constraint, we introduce DeciMamba, a context-extension method specifically designed for Mamba. This mechanism, built on top of a hidden filtering mechanism embedded within the S6 layer, enables the trained model to extrapolate well even without additional training. Empirical experiments over real-world long-range NLP tasks show that DeciMamba can extrapolate to context lengths that are 25x times longer than the ones seen during training, and does so without utilizing additional computational resources. We will release our code and models.
https://github.com/assafbk/DeciMamba
@@
Assaf Ben-Kish1   Itamar Zimerman1   Shady Abu-Hussein1 Nadav Cohen1   Amir Globerson1,2   Lior Wolf1   Raja Giryes1 1Tel Aviv University     2Google Research
@@ 1 Introduction
Figure 1: Document Retrieval Over SQuAD. We present a novel decimation mechanism tailored for Mamba, enabling it to process sequences that are x25 times longer than the ones seen during training. The purple line shows the training sequences length. Results are shown for DeciMamba-130M.
@@
Lengthy sequences, which can span up to millions of tokens, are common in real-world applications including long books, high-resolution video and audio signals, and genomic data. Consequently, developing Deep Learning (DL) sequence models capable of effectively managing long contexts is a critical objective. Transformers Vaswani et al. (2017), despite their current dominance in general DL tasks, still face challenges in processing long sequences. Specifically, their quadratic complexity in sequence length makes them computationally demanding, restricting the ability to train them over long sequences and very large datasets.
@@
In recent years, substantial efforts have been made in order to tackle this challenge. The most significant advancements include efficient implementations that increase the model’s context length during training Dao et al. (2022); Liu et al. (2023), and context-extension methods Chen et al. (2023b); Peng et al. (2023b) designed to effectively expand the context after training. However, recent studies suggest that long-range processing is still an unresolved problem Li et al. (2024); Liu et al. (2024a).
@@
One promising approach in this domain is the development of attention-free networks with sub-quadratic complexity, which can be trained more effectively over long sequences. In a recent line of work Gu et al. (2021b, a); Gupta et al. (2022), the family of state-space layers has been introduced. These layers can be seen as theoretically grounded linear RNNs that can be efficiently computed in parallel via convolutions, thanks to a closed-form formulation of their linear recurrent rule. A recent advancement by Gu and Dao (2023) presented Mamba, which builds on top of an expressive variant of SSMs called Selective State-Space Layers (S6). These layers match or exceed the performance of Transformers in several domains, such as NLP Pióro et al. (2024); Wang et al. (2024), image classification Zhu et al. (2024); Liu et al. (2024b), audio processing Shams et al. (2024), genomic data Schiff et al. (2024), and more.
@@
In this paper, we first explore the length-generalization abilities of Mamba and identify that they are relatively limited. Although Mamba layers are theoretically capable of capturing global interactions at the layer level, through a series of visualizations, analyses, and empirical measures, we show that the main barrier preventing Mamba from extrapolating is its implicit bias towards a limited effective receptive field (ERF), which is bounded primarily by the sequence length used during training. Secondly, based on the assumption that long-context data is usually sparse, we present DeciMamba, the first context-extension method specifically designed for S6. Our method relies on a dynamic data-dependent pooling method that utilizes a hidden filtering mechanism intrinsic to the Mamba layer. We leverage this mechanism to introduce a global compression operator, which expands Mamba’s ERF by discarding un-important tokens before each S6 layer. Our method significantly increases the effective context length of Mamba by several orders of magnitude while requiring a smaller computational budget.
@@
Our main contributions encompass the following three aspects: (i) identify that Mamba has limited length-extrapolation capabilities, (ii) recognize via a series of visualizations, analyses, and empirical measures that although Mamba can theoretically capture global interactions via the recurrent state, its limited ERF prevents significant length-extrapolation. (iii) Building on this insight, we introduce DeciMamba, the first context-extension technique specifically designed for Mamba models. This approach leverages an existing filtering mechanism embedded within the S6 layer. As illustrated in Fig. 9, our method effectively enhances Mamba’s length-extrapolation abilities by up to eightfold, and is applicable to real-world long-context NLP tasks, as demonstrated in Fig. 1.
@@ 2 Preliminaries
In this section, the scientific context for discussion of long-range models is described and the necessary terminology and symbols are provided. Such models evolve in two main directions: (i) first, by adapting transformers, the most dominant architecture today, to be more suitable for such tasks, or, alternatively, (ii) by developing architectures with sub-quadratic complexity in sequence length. Several modern examples include Hyena Poli et al. (2023), RWKV Peng et al. (2023a), Hawk De et al. (2024), xLSTM Beck et al. (2024), and Mamba, the latter being the focus of our paper.
@@
$G=\mathrm{\sigma}(\text{Linear}(U)),\quad X=\text{Conv1D}(\text{Linear}(U)),$ (1)
$Y=S6(X),\quad O=Y\otimes G$ (2)
@@
$\Delta_{t}=\textit{Sft}(S_{\Delta}X_{t}),\quad B_{t}=S_{B}X_{t},\quad C_{t}=(S_{C}X_{t})^{T}$ (4)
@@
such that $\Delta_{t}$ is the discretization step, Sft represents the softplus function, and $S_{\Delta},S_{B},S_{C}$ are linear projection layers. Ali et al. (2024) demonstrated that S6 layers, similar to attention models, can be interpreted as data-controlled linear operators. Specifically, the S6 layer computation can be represented using the following linear operator $\alpha$, controlled by the input (via Eq. 4 and 5):
@@
Context Extension & Length Extrapolation. Over the years, several methods have been proposed to enhance the effective context length of transformers and improve their ability to extrapolate over longer sequences. Traditionally, length generalization is closely related to positional encoding. Pioneering work in the domain, introduced by Press et al. (2021), demonstrates that models built on top of original sinusoidal, rotary Su et al. (2024), and T5 bias Raffel et al. (2020) positional encoding have poor length generalization abilities. It proposes Alibi, which mitigates this issue by promoting locality through the incorporation of distance-based linear biases into the attention matrix. A counter-intuitive study was conducted by Kazemnejad et al. (2024), which shows that transformers without positional encoding (NoPE) exhibit better length extrapolation capabilities in downstream tasks. Two more promising approaches are the very recent CoPE Golovneva et al. (2024), which utilizes context-aware positional encoding, and post-training positional interpolation Peng et al. (2023b); Chen et al. (2023a).
Alternatively, a recent direction involves architectural modifications to pre-trained models followed by short fine-tuning. Such works include LongLora Chen et al. (2023b), which proposes shifted sparse attention, and Landmark Attention Mohtashami and Jaggi (2023), which applies attention in chunks and inserts global unique tokens into the input sequences between those chunks. Our method can be considered an application of this approach to Mamba models, rather than to traditional transformers. More related work is in App. B.
We note that in order to assess the long range performance of the model, in our experiments we measure perplexity only on the farthest labels of the context window, causing the perplexity values to be larger than their typical values (usually, all labels in the context window are aggregated).
@@ 3 Extrapolation Limitations of Mamba
In this section we explore the length-generalization capabilities of Mamba models. Specifically, we assess the effective receptive field (ERF) of S6 layers, a phenomenon that leads to poor information propagation when processing sequences that are longer than the ones trained on.
To investigate why Mamba fails at long sequence extrapolation we visualize Mamba’s hidden attention (Eq. 67). In Fig. 2 we display the attention matrices of Mamba-130M after training it on the Passkey Retrieval task (Sec. 5) with a sequence length of 2K. We record the data for the matrices in layer 17, while evaluating sequences of length 2K and 16K. Notice that while the attention maps are very dense for the 2K sequence inference, the attention maps for the 16K sequence are much more sparse, and their elements vanish as we move towards the lower rows. This exhibits a limited effective receptive field (ERF), as information from the first tokens in the sequence does not propagate to the final tokens in the output sequence.
Figure 2: ERF Detected by the Mamba Attention Matrix (Eq. 7). We feed Mamba-130m with two sequences of different lengths, 2K and 16K, and show the attention matrices of layer 17. Mamba unintentionally learns an effective receptive field during training (right image, highlighted by the dashed rectangle), a behavior which limits its extrapolation abilities.
Measuring ERFs via Mamba Mean Distance. To understand the trend observed in Fig. 2, we introduce a quantitative measure called ‘Mamba Mean Distance’, which estimates the ERF of Mamba layers. This measurement is analogous to the receptive field in CNNs and the attention mean distance used in transformers, as described in the seminal work of Dosovitskiy et al. (2020). For a causal transformer model, the attention mean distance for the $i$-th output token is computed as follows:
@@
Where $\tilde{A}$ is the normalized attention matrix that defines a probability distribution over the distances for various tokens. We tailored this mechanism to S6 by leveraging the implicit attention representation, which we normalize using the following normalization function:
$\mathbb{N}(x_{1},\cdots,x_{L})_{j}=\frac{|x_{j}|}{\sum_{i=1}^{L}|x_{i}|}.$ (9)
Since we are usually interested in the last tokens, and as we will see shortly, they are the first to suffer from this phenomenon, we find that it is sufficient to compute our metric for the last token only:
$\vspace{-3pt}\mathop{\mathbb{E}}_{j\leq L}d(j,L)\approx\sum_{j\leq L}\mathbb{N}(\alpha)_{L,j}\otimes(L-j).\vspace{-3pt}$ (10)
Fig. 4 highlights the significance of this measure by presenting the implicit attention matrices of Mamba layers alongside their corresponding ‘Mamba Mean Distance’, which is depicted by the distance of the diagonal red line from the main diagonal of the attention matrix.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3: Mamba Mean Distance. Each panel contains an attention matrix along with its corresponding ’Mamba Mean Distance’, depicted by the distance of the red diagonal line from the main diagonal. The matrices are extracted from a pre-trained model of size 2.8B, trained on the Pile dataset Gao et al. (2020).
@@
Figure 4: Mamba Mean Distance Quantifies Extrapolation Abilities. (Left) Mamba Mean Distance as a function of the context length during inference, for various training context lengths. (Right) Same, but normalized by the training context length.
@@
By introducing an empirical measurement for Mamba’s ERF, we can explore it more robustly. Fig. 4 depicts how the Mamba Mean Distance is influenced when increasing the input sequence length during training and inference. Here, Mamba-based language models with 80M parameters are trained across various context lengths for next-token prediction on the WikiText-103 benchmark Merity et al. (2016). We average the Mamba Mean Distance over all channels and layers using 100 test examples.
The left panel of Fig. 4 shows that the Mamba Mean Distance increases with longer context lengths during both training and evaluation. However, this does not take into account the length of the context that needs attending to. In the right panel, we normalize the Mamba Mean Distance by the sequence length during inference, revealing that the proportion of context utilized in practice decreases as the context length during evaluation increases. This phenomenon suggests that the limited length-extrapolation capabilities of Mamba models are due to a reduced ERF (represented by the normalized Mamba Mean Distance), which diminishes as the context length increases during evaluation.
@@
To better understand this phenomenon, we return to the implicit attention defined in Eq. 7. We note that the product of transition matrices $\prod_{k=j+1}^{i}\bar{A}_{k}$ includes more elements as the sequence length $L$ increases, and can be written for the last row as:
$\vspace{-3pt}\prod_{k=j+1}^{L}\bar{A}_{k}=\prod_{k=j+1}^{L}\exp(A\Delta_{k})=\exp(A\sum_{k=j+1}^{L}\Delta_{k})$ (11)
@@
We note that this product always converges since $\Delta_{t}\geq 0$ and $A[i,i]<0$ by design. Yet, if it converges too fast, the attention element will collapse undesirably. To quantify the loss of information we focus on the attention element $\alpha_{L,1}$ due to the following reasons: (i) we are usually interested in the information at the end of the sequence; (ii) this element is the first in the row to collapse, capturing the first occasion of information loss because $\sum_{k=2}^{L}\Delta_{k}\geq\sum_{k=j+1}^{L}\Delta_{k}$ for all $L-1\geq j\geq 1$, and $A$ is always negative, so the power of its exponent will always be the smallest. We measure $\sum_{k=2}^{L}\Delta_{k}$ in Fig. 6 for varying evaluation context lengths L, per layer, averaged over the channels dimension. As can be seen in the figure, the values of $\sum_{k=2}^{L}\Delta_{k}$ increase exponentially fast in the most semantic layers (layers 16 and 17, see Passkey Retrieval in Sec. 5 for more details), leading to a fast collapse. Note that these are the layers where the worst ERFs are observed. The values of $\sum_{k=2}^{L}\Delta_{k}$ grow linearly in the other layers, which perform local processing (Fig. 15). Their decay is intentional.
A likely explanation for the limited length-generalization abilities is that during training, each SSM layer learns decay rates $A$ that are sufficient for processing training data of length $L_{train}$ but not flexible enough to provide effective length-extrapolation for lengths $L_{eval}>L_{train}$ Specifically, if the decay rate is too high, long dependencies cannot be captured. Conversely, if the decay rate is too low, the state becomes “overloaded” with excessive information, which can obscure the important signals needed for effective processing.
Figure 5: Quantifying Mamba’s Information Loss by Measuring $\sum_{k=2}^{L}\Delta_{k}$ Divergence. To show Mamba’s sensitivity to increasing context lengths we measure the first occasion of information loss, as described in Sec.3. We observe that in the most semantic layers (16 and 17, see Passkey Retrieval in Sec. 5) $\sum_{k=2}^{L}\Delta_{k}$ diverges exponentially fast, causing a fast decay in its respective attention value, leading to the ERF observed in Fig. 2.
@@
Figure 6: Perplexity Over PG-19 Test Set. The dashed purple line shows the train sequence length for Mamba and DeciMamba (Mamba-130m). Each point on the green curve shows a different Mamba model trained on the respective context length ($L_{train}=L_{eval}$); for $L_{train}>30K$ an Out Of Memory (OOM) error occurs.

### v2

@@ Abstract
Long-range sequence processing poses a significant challenge for Transformers due to their quadratic complexity in input length. A promising alternative is Mamba, which demonstrates high performance and achieves Transformer-level capabilities while requiring substantially fewer computational resources. In this paper we explore the length-generalization capabilities of Mamba, which we find to be relatively limited. Through a series of visualizations and analyses we identify that the limitations arise from a restricted effective receptive field, dictated by the sequence length used during training. To address this constraint, we introduce DeciMamba, a context-extension method specifically designed for Mamba. This mechanism, built on top of a hidden filtering mechanism embedded within the S6 layer, enables the trained model to extrapolate well even without additional training. Empirical experiments over real-world long-range NLP tasks show that DeciMamba can extrapolate to context lengths that are significantly longer than the ones seen during training, while enjoying faster inference.
@@
Anonymous ACL submission
@@ 1 Introduction
Lengthy sequences, which can span up to millions of tokens, are common in real-world applications including long books, high-resolution video and audio signals, and genomic data. Consequently, developing Deep Learning (DL) sequence models capable of effectively managing long contexts is a critical objective. Transformers (vaswani2017attention), despite their current dominance in general DL tasks, still face challenges in processing long sequences. Specifically, their quadratic complexity in sequence length makes them computationally demanding, restricting the ability to train them over long sequences and very large datasets.
@@
In recent years, substantial efforts have been made in order to tackle this challenge. The most significant advancements include efficient implementations that increase the model’s context length during training (dao2022flashattention; liu2023ring), and context-extension methods (chen2023longlora; peng2023yarn) designed to effectively expand the context after training. However, recent studies suggest that long-range processing is still an unresolved problem (li2024long; liu2024lost).
Figure 1: Improving Mamba Extrapolation with DeciMamba. We present a novel decimation mechanism tailored for Mamba. With our method Mamba is able to process sequences that are significantly longer than the ones trained on while enjoying reduced inference costs. (Left) Document Retrieval. (Right) Passkey Retrieval. In all plots, the purple line shows the training sequence length. Results are shown for DeciMamba-130M.
@@
One promising approach in this domain is the development of attention-free networks with sub-quadratic complexity, which can be trained more efficiently over long sequence data. In a recent line of works (gu2021combining; gu2021efficiently; gupta2022diagonal), the family of state-space layers has been introduced. These layers can be seen as theoretically grounded linear RNNs that can be efficiently computed in parallel via convolutions, thanks to a closed-form formulation of their linear recurrent rule. A recent advancement by gu2023mamba presented Mamba, which builds on top of an expressive variant of SSMs called Selective State-Space Layers (S6). These layers match or exceed the performance of Transformers in several domains, such as NLP (lieber2024jamba; zuo2024falcon; pioro2024moe; wang2024mambabyte), image classification (zhu2024vision; liu2024vmamba), audio processing (shams2024ssamba), genomic data (schiff2024caduceus), and more.
@@
In this paper, we first explore the length-generalization abilities of Mamba and identify that they are relatively limited. Although Mamba layers are theoretically capable of capturing global interactions at the layer level, we show, through a series of visualizations, analyses, and empirical measures, that the main barrier is Mamba’s implicit bias towards sequence lengths that were seen during training, a phenomenon that we call ‘limited effective receptive field’ (ERF). Next, based on the assumption that long-context data is usually sparse, we present DeciMamba, the first context-extension method specifically designed for S6. Our method relies on a dynamic data-dependent pooling method that utilizes a hidden filtering mechanism intrinsic to the Mamba layer. We leverage this mechanism to introduce a global compression operator, which expands Mamba’s ERF by discarding unimportant tokens before the S6 layer. In particular, we interpret the norms of the per-token selective recurrent gate ($\Delta_{t}$) as indicators of each token’s importance. This metric allows us to identify the top-k most impactful tokens, enabling direct application of the SSM to these tokens only. The proposed method (Figure 2) significantly increases the effective context length of Mamba by several orders of magnitude while requiring a smaller computational budget.
@@
Our main contributions encompass the following three aspects: (i) identify that Mamba has limited length-extrapolation capabilities via thorough experiments in controlled environments. (ii) Through a series of visualizations, analyses, and empirical measures, recognize that although Mamba can theoretically capture global interactions via the recurrent state, its limited ERF prevents significant length-extrapolation; (iii) building on this insight, introduce DeciMamba, the first context-extension technique specifically designed for Mamba models. This approach leverages an existing filtering mechanism embedded within the S6 layer. As illustrated in Fig. 1, our method effectively enhances Mamba’s length-extrapolation abilities, and is applicable to real-world long-context NLP tasks.
Figure 2: DeciMamba (Left) Schematic overview; (Middle) By carefully inspecting the recurrent view of Mamba, we revealed the implicit filtering mechanism embedded in the recurrent gate and controlled by $\Delta_{t}$; (Right) Visualization of a DeciMamba model. The grey lines represent the sequence length at the input and output of each layer. Layers with a large ERF are decimated.
@@ 2 Preliminaries
Long-range models evolve in two main directions: (i) adapting transformers, the most dominant architecture today, to be more suitable for such tasks; (ii) developing architectures with sub-quadratic complexity in sequence length such as Hyena (poli2023hyena), RWKV (peng2023rwkv), Hawk (de2024griffin), xLSTM (beck2024xlstm), and Mamba, the focus of our paper.
@@
$G=\mathrm{\sigma}(\text{Linear}(U)),$ $X=\text{Conv1D}(\text{Linear}(U)),$ (1)
$Y=S6(X),$ $O=Y\otimes G$
@@
$\Delta_{t}=\textit{Sft}(S_{\Delta}X_{t}),B_{t}=S_{B}X_{t},C_{t}=(S_{C}X_{t})^{T}$
@@
such that $\Delta_{t}$ is the discretization step, Sft represents the softplus function, and $S_{\Delta},S_{B},S_{C}$ are linear projection layers. ali2024hidden demonstrated that S6 layers, similar to attention models, can be interpreted as data-controlled linear operators. Specifically, the S6 layer computation can be represented using the following linear operator $\alpha$, controlled by the input (via Eq. 2):
@@
Context Extension & Length Extrapolation. Several methods were proposed to enhance the effective context length of transformers and improve their extrapolation over longer sequences. press2021train demonstrated that models built on top of original sinusoidal, rotary (su2024roformer), and T5 bias (raffel2020exploring) positional encoding have poor length generalization. It proposed to mitigate this issue by incorporating distance-based linear biases into the attention matrix for promoting locality . kazemnejad2024impact showed that transformers without positional encoding (NoPE) exhibit better length extrapolation capabilities in downstream tasks. Recently, CoPE (golovneva2024contextual) utilized context-aware positional encoding and peng2023yarn; chen2023extending suggested post-training positional interpolation.
A recent direction involves architectural modifications to pre-trained models followed by short fine-tuning. It includes LongLora (chen2023longlora), which proposes shifted sparse attention, and Landmark Attention (mohtashami2023landmark), which applies attention in chunks and inserts global unique tokens into the input sequences between those chunks. Our work focuses on taking such an approach for Mamba models, rather than transformers. More related work is in Appendix D. Furthermore, we refer to Appendix D.4, where we motivate the study of Long Context LLMs in parallel to Retrieval Augmented Generation (RAG). Lastly, we note that in order to assess the long range performance of the model, in our experiments we measure perplexity only on the farthest labels of the context window, causing the perplexity values to be larger than their typical values (usually, all tokens in the context window are aggregated).
@@ 3 Extrapolation Limitations of Mamba
In this section we explore the length-generalization capabilities of Mamba models. Specifically, we assess the limited effective receptive field (ERF) of S6 layers, a phenomenon that leads to poor information propagation when processing sequences that are longer than the ones trained on. For a detailed description of the ERF, we refer the reader to Appendix D.3. Consequently, we develop methods to measure the portion of the context utilized by the model, and show that it is dictated by the training sequence lengths.
We start by visualizing Mamba’s hidden attention (Eq. 45) during extrapolation. In Fig. 3 (left, center) we display the attention matrices of layer 17 of Mamba-130M while evaluating sequences of length 2K and 16K (the model trained on the Passkey Retrieval task with sequence lengths of 2K). Notice that while the attention is very dense for the 2K sequence (no extrapolation), the attention for the 16K sequence (extrapolation x8) is much more sparse, and its elements vanish as we move towards the lower rows. This exhibits a limited ERF, as information from the first tokens in the sequence does not propagate to the final tokens in the output sequence.
Figure 3: Detecting and Quantifying Limited ERFs. (Center, Left) Recordings of Mamba Attention Matrices with and without extrapolation (Mamba-130m, layer 17, trained on seq. lengths of 2k). Mamba unintentionally learns a limited ERF during training (highlighted by the dashed rectangle) which disrupts its extrapolation abilities. (Right) Quantifying Mamba’s Information Loss by Measuring $\sum_{k=2}^{L}\Delta_{k}$ Divergence. To show Mamba’s sensitivity to increasing context lengths we measure the first occasion of information loss, as described in Sec. 3. We observe that in the most semantic layers (16 and 17, see Passkey Retrieval in Sec. 2) $\sum_{k=2}^{L}\Delta_{k}$ diverges exponentially fast, causing a fast decay in the attention values, leading to limited ERFs like in the center image.
Measuring ERFs via Mamba Mean Distance. To quantify how well Mamba utilizes the context during inference, we introduce a quantitative measure called ‘Mamba Mean Distance’. This measure is analogous to the receptive field in CNNs and the attention mean distance in transformers described by dosovitskiy2020image. For a causal transformer model, the attention mean distance for the $i$-th output token is computed by:
@@
where $\tilde{A}$ is the normalized attention matrix that defines a probability distribution over the distances for various tokens. We tailored this measure to S6 by leveraging the implicit attention representation, which we normalize using the following function:
$\mathbb{N}(x_{1},\cdots,x_{L})_{j}=\frac{|x_{j}|}{\sum_{k=1}^{L}|x_{k}|}.$ (7)
As we are usually interested in the last tokens, and as we will see shortly, they are the first to suffer from this phenomenon, we find that it is sufficient to compute our metric for the last token only:
$\mathop{\mathbb{E}}_{j\leq L}d(L,j)\approx\sum_{j\leq L}\mathbb{N}(\alpha)_{L,j}\otimes(L-j).$ (8)
Fig. 4 visualizes this measure by presenting the ‘Mamba Mean Distance’ for different hidden attention matrices, depicted by the horizontal distance between the red line and the main diagonal.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Mamba Mean Distance. Each panel contains an attention matrix along with its corresponding ’Mamba Mean Distance’, depicted by the horizontal distance between the red diagonal line and the main diagonal. The matrices are extracted from a pre-trained model of size 2.8B, trained on the Pile dataset (gao2020pile).
@@
Figure 5: Mamba Mean Distance Quantifies Effective Context Utilization. (Left) Mamba Mean Distance as a function of the context length during inference, for various training context lengths. (Right) Same, but normalized by the training context length.
@@
By introducing an empirical measure for Mamba’s ERF, we can explore it more robustly. Fig. 5 portrays the relationship between the input context length and the Mamba Mean Distance. Here, Mamba-based language models with 80M parameters are trained across various context lengths for next-token prediction on the WikiText-103 benchmark (merity2016pointer). We average the Mamba Mean Distance over all channels and layers using 100 test examples. More details on Mamba Mean Distance are in Appendix D.5. Surprisingly, in the left panel of Fig. 5, we see that the Mamba Mean Distance increases with context length. Although it does grow, it is hard to tell whether the growth is proportional to the growth in context length. Therfore, in the right panel, we normalize the Mamba Mean Distance by the sequence length during inference, revealing the exact proportion of context utilized in practice. It is now evident that the context utilization drops dramatically as the evaluated context length increases.
@@
To identify the internal mechanism that leads to this behavior, we return to the analytic expression of Mamba’s hidden attention defined in Eq. 5. We observe that the product of transition matrices $\prod_{k=j+1}^{i}\bar{A}_{k}$ includes more elements as the sequence length $L$ increases, and can be written for the j’th element in the last row as:
$\prod_{k=j+1}^{L}\bar{A}_{k}=\prod_{k=j+1}^{L}\exp(A\Delta_{k})=\exp(A\sum_{k=j+1}^{L}\Delta_{k})$ (9)
@@
We note that this product always converges since $\Delta_{t}\geq 0$ and $A[i,i]<0$ by design. Yet, if it converges too fast, the attention element will collapse undesirably. To measure the attention elements’ collapse we focus on attention element $\alpha_{L,1}$ due to the following reasons: (i) we are usually interested in the information at the end of the sequence; (ii) this element is the first in the row to collapse, capturing the first occasion of information loss because $\sum_{k=2}^{L}\Delta_{k}\geq\sum_{k=j+1}^{L}\Delta_{k}$ for all $L-1\geq j\geq 1$, and $A$ is always negative, so the power of its exponent will always be the smallest. As can be seen in Eq. 9, the only input-length dependent value is $\sum_{k=2}^{L}\Delta_{k}$. Therefore, in Fig. 3 (right), we compute it for varying input lengths L during inference (per layer, averaged over the channels dimension). As can be seen in the figure, the values of $\sum_{k=2}^{L}\Delta_{k}$ start small and increase exponentially fast in layers 16 and 17, which are the most global layers (Fig. 13). This aligns with the observations in Fig.  3 (left, center) - since these layers require a large ERF, the fast collapse of the attention elements restricts global information propagation, leading to distorted performance. In the other layers, the values of $\sum_{k=2}^{L}\Delta_{k}$ start large and grow linearly. Since these layers perform local processing (diagonal attention matrices), the fast decay is desired.
To conclude, the limited length-generalization abilities arise due to decay rates $A$ and $\Delta_{t}$ sums (Eq. 9) that are sufficient for the training data of length $L_{train}$, but not flexible enough to provide effective length-extrapolation for lengths $L_{eval}>L_{train}$. As the transition matrix product converges too fast, we observe a collapse of the attention matrix, specifically in regions that convey information from the beginning of the sequence to its end.
@@
Figure 6: Perplexity Over PG-19 Test Set. (Left) The dashed purple line shows the train sequence length for Mamba-130m and DeciMamba-130m. Each point on the green curve shows a different Mamba-130m model trained on the respective context length ($L_{train}=L_{eval}$); for $L_{train}>30K$ an Out Of Memory (OOM) error occurs. (Right) Zero-Shot perplexity comparison.